OpenStudy (anonymous):

Find an equation of the line that passes through the given points. (Let x be the independent variable and y be the dependent variable.) (4, 14) and (5, 19)

OpenStudy (anonymous):

Ok, first off, you are going to have to find the slope.

OpenStudy (anonymous):

and i do that how? :D

OpenStudy (anonymous):

m = (y2-y1) / (x2-x1)

OpenStudy (anonymous):

So, you are going to set the problem up like this... m = (19-14)/(5-4)

OpenStudy (anonymous):

m = 5/1 m = 5

OpenStudy (anonymous):

Next, you will use the equation y-y1=m(x-x1)

OpenStudy (anonymous):

why that equation?

OpenStudy (anonymous):

y-14=5(x-4)

OpenStudy (anonymous):

Because it gives you an equation answer as you ask. This kind of equation is known as "point-slope" form.

OpenStudy (anonymous):

Your final answer... y-14=5x-20 y=5x+34